# HTMformer: Hybrid Time and Multivariate Transformer for Time Series Forecasting

### v2

@@ Appendix G Details of Model Efficiency
For a comprehensive efficiency comparison, we evaluate HTMformer against three highly competitive baselines, including MultiPatchFormer (Naghashi et al., 2025), WPMixer (Murad et al., 2025), and PatchTST (Nie et al., 2023), across eight datasets. The evaluation considers three key metrics: training time, GPU memory footprint, and total parameter count, with the input sequence length fixed at 96 and the prediction length at 192. Moreover, the experimental hardware conditions and parameter configurations follow those described in Appendix B, and the best results are marked in red. The full results are shown in Table 14 (Left) and Table 14 (Right). The training time and GPU usage serve as an intuitive indicator of total parameter count, so detailed reporting is omitted. Across all datasets, HTMformer achieves the best training time and GPU memory footprint. Notably, for high-dimensional datasets, particularly when the number of variate dimensions exceeds one hundred, HTMformer delivers nearly three times the training speed of the second-best model and requires only one-third of its memory footprint. Moreover, this advantage becomes increasingly pronounced as dimensionality grows. This superiority stems from our lightweight design philosophy, in which HTME avoids overemphasizing temporal feature extraction and HTMformer adopts only a vanilla Transformer encoder combined with the inverted input strategy.
@@
Models Ours 2025 2025 2023
@@
Models Ours 2025 2025 2023
@@
Appendix H Hyperparameter Sensitivity
@@
Appendix I Visualization of Prediction Results
We visualize HTMformer’s ability to predict trends across various time series datasets, including ECL, Traffic, Solar-Energy, Weather, and PEMS (PEMS03, PEMS04 and PEMS08), as shown in Figures˜7, 8, 9, 10, 11, 12, 13 and 14. Each example uses a 96-step input to generate 96-step predictions. In the visualizations, the orange lines indicate the ground truth values, and the blue lines show the model’s predictions. It accurately captures the cyclical patterns and oscillatory behaviors, and successfully forecasts the overall directional trends.
@@ Appendix I Visualization of Prediction Results
Appendix J Discussion on Transformer-based Predictors
Recent studies (Lu et al., 2023; Li et al., 2023b) have found that Channel Dependence ideally gains from higher capacity, while Channel Independence can significantly enhance performance due to sample scarcity. In time series forecasting, an increasing number of models employ Channel Independence (Nie et al., 2023), as most current forecasting benchmarks are not sufficiently large. These models treat time series variates independently and utilize a shared backbone, achieving good performance. However, Channel Independence inevitably overlooks the multivariate correlations that exist in most datasets, resulting in the loss of some information. We believe that making variates independent is desirable, but multivariate correlations should not be ignored. However, previous improvements to models have mainly focused on extracting features in the temporal dimension. Excessive extraction of temporal features does not further improve prediction accuracy, but rather significantly increases computational overhead.
Notably, as demonstrated in both the introduction and ablation studies of this paper. The main reason why traditional Transformer-based predictors encounter performance bottlenecks is that the embedding model extracts only limited effective information from the sequence, which in turn affects the performance of subsequent models. Due to the inherent limitations of Transformer-based predictors, simply enlarging the input sequence length cannot further enhance accuracy. Constructing an embedding model that deeply extracts sequence information can effectively address this problem. Therefore, we propose a novel strategy, specifically extracting multivariate features in the embedding layer and using them to complement temporal features. This approach incorporates multivariate correlations, thereby further extracting effective information from the sequence. Moreover, it does not affect the subsequent attention mechanism to utilize Channel Independence.
We find that extracting local patterns in time series through patching and convolution, together with capturing long-term dependencies using linear layers, is sufficient for effectively modeling temporal features. For Transformer-based predictors, this is sufficient to extract temporal representations, and it is unnecessary to employ attention mechanisms to model temporal correlations.
Graph Convolutional Networks (GCNs) typically construct an adjacency matrix to represent multivariate correlations (Kipf & Welling, 2017; Bruna et al., 2014). However, the adjacency matrix is often non-learnable. Although the physical locations of sensors are relatively fixed, the correlations among variables are not necessarily constant, which severely affects the performance of GCNs. Introducing a learnable adjacency matrix (Wang et al., 2019) can substantially increase computational overhead and lead to model instability (Wu et al., 2020a; Franceschi et al., 2019). Therefore, this paper adopts a "weak learning" strategy, whereby multivariate correlations are simply extracted and incorporated as features into tokens to correct temporal features.
We validate the effectiveness and scalability of the proposed HTME strategy by integrating it into various Transformer variants. HTME can be seamlessly combined with various attention mechanisms across a wide spectrum of time series forecasting tasks to achieve optimal performance. Our model demonstrates superior performance in short-term sequence forecasting. On the selected datasets, PEMS03, PEMS04, and PEMS08, our model achieves the best results compared with the second-best methods. In long-term sequence forecasting, our model ranks first or second on eight selected datasets in most cases. This demonstrates the effectiveness of the HTME strategy in advancing time series forecasting. We also compared the Inference Time, GPU Memory Footprint, and Model Parameters of HTMformer with those of three models that have good forecasting performance: MultiPatchFormer, WPMixer, and PatchTST. On the selected datasets, HTMformer achieves superior or equivalent performance with only about 0.2-0.45 times the memory of the three models, and is two to three times faster. These results suggest that the HTME model is highly flexible for deployment across a broader spectrum of devices and can better accommodate real-time forecasting tasks, thereby creating greater potential for practical applications and future scalability.

### v3

@@ Appendix G Details of Model Efficiency
For a comprehensive efficiency comparison, we evaluate HTMformer against three highly competitive baselines, including MultiPatchFormer (naghashi2025multipatchformer), WPMixer (murad2024wp), and PatchTST (nie2023time), across eight datasets. The evaluation considers three key metrics: training time, GPU memory footprint, with the input sequence length fixed at 96 and the prediction length at 192. Moreover, the experimental hardware conditions and parameter configurations follow those described in Appendix D, and the best results are marked in red. The full results are shown in Table 13 (Left) and Table 13 (Right). Across all datasets, HTMformer achieves the best training time and GPU memory footprint. Notably, for high-dimensional datasets, particularly when the number of variate dimensions exceeds one hundred, HTMformer delivers nearly three times the training speed of the second-best model and requires only one-third of its memory footprint. Moreover, this advantage becomes increasingly pronounced as dimensionality grows. HTME is a lightweight module that enables simple models focused on temporal feature modeling to incorporate the influence of multivariate correlations, thereby achieving performance comparable to that of complex models.
@@
Models Ours naghashi2025multipatchformer murad2024wp nie2023time
@@
Models Ours naghashi2025multipatchformer murad2024wp nie2023time
@@
Appendix H Analysis of HTME
The vast majority of features in time series are contained within the temporal dimension, and the depth of exploration in this dimension significantly influences prediction accuracy (hyndman2021forecasting). Therefore, existing models can achieve superior performance even by focusing solely on the temporal dimension. Moreover, thanks to their simple model architectures, they often have a low computational overhead.
The multivariate dimension also contains a large number of features (Pravilovic2013SpatialForecasting), and fully exploiting multivariate features can further boost performance. For time series data, the vast majority of features are latent within the temporal dimension, making the deep extraction of temporal features the key to effective modeling. Overemphasizing multivariate features—particularly strategies that explicitly model cross-variable correlations within the backbone network—introduces significant computational overhead without a corresponding performance gain. Most importantly, introducing additional complex modules into the backbone network to model multivariate correlations forces temporal and multivariate features to share the same feature space. Consequently, while the model is capturing cross-variable relationships, the representation of temporal features is inevitably compromised.
We visualize the temporal and multivariate feature spaces of HTME, as well as the weight parameter alpha. The value of alpha is approximately 0.65, indicating that subsequent modules are more inclined to select temporal features. Additionally, temporal feature weights are considerably higher than spatial feature weights. This emphasizes the primary role of temporal information.
We also observe that the temporal features exhibit a strip-like distribution, showing significant variation along the temporal dimension but minimal differences in the multivariate dimension. This suggests that the temporal features possess a high degree of consistency across the multivariate dimension, which in turn underscores their importance. The multivariate features exhibit a point-like and uniform distribution. For the most part, they take on small values, which ensures that they do not disrupt the temporal feature space. A few points within the multivariate features exhibit large values. These points serve to emphasize their corresponding temporal features, thereby creating a more information-rich temporal feature space. Consequently, subsequent modules can utilize this enhanced feature space more efficiently, improving the model’s capability to capture complex temporal dynamics.
Appendix I Hyperparameter Sensitivity
@@
Appendix J Visualization of Prediction Results
We visualize HTMformer’s ability to predict trends across various time series datasets, including ECL, Traffic, Solar-Energy, Weather, and PEMS (PEMS03, PEMS04 and PEMS08), as shown in Figures˜8, 9, 10, 11, 12, 13, 14 and 15. Each example uses a 96-step input to generate 96-step predictions. In the visualizations, the orange lines indicate the ground truth values, and the blue lines show the model’s predictions. It accurately captures the cyclical patterns and oscillatory behaviors, and successfully forecasts the overall directional trends.